News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 12 June 2000 8.30am EDT

Last week was an eventful week for the Chandra Team.

1. Command Timing Overlap

During the 8am EDT pass on 6/5 the spacecraft was not in the configuration expected from the daily load. The reason was traced to a timing overlap between the completion of one command and the start of the next. The first command (which issued a set of ACIS commands) was completed after the absolute start time of the next command (an HRC door open command), and resulted in the OBC waiting in a loop for the correct time to occur - effectively waiting for many days. The event occurred at 3:02am EDT on 6/5. After assessing nominal spacecraft health, SCS 107 was run in real-time to stop the loads and safe the Science Instruments. Real-time procedures were run to re-enable momentum monitors, to enable momentum auto-unloading, and to re-open the HRC door to the partially closed position. The daily loads were rebuilt to continue science operations beginning at 11:59pm EDT on 6/5. A series of calibration observations of PSR B0540-69 were impacted by the event, and will be rescheduled in an upcoming load. Checks have been implemented to detect and avoid similar timing overlaps in future loads, and modifications are being assessed to the ground software.

2. High Radiation Trigger

At 5:41 EDT on 6/8 the ACIS high radiation flag in the Flight Software Radiation Monitor was set to TRUE, resulting in activation of SCS 107 to safe the science instruments. This was due to high counts for protons in the EPHIN P4GM channel, exceeding the threshold of 300 counts/cm^2/s/sr (the channel reached 460). The high proton counts occurred following passage of an interplanetary shock wave due to a halo coronal mass ejection associated with two X-class solar flares on 6 June. SCS 107 ran nominally and all safing actions were completed with the exception of the HETG retraction. The HETG was not retracted because the MCE-B patch has not yet been uplinked - this is planned for next week. This was the first occurrence of an SI safing action due to high radiation environment for the mission.

The observations impacted by the trigger were MCG-6-30-15 and X0918-54 which will be rescheduled in an upcoming load. The daily load was rebuilt to restart at 8.30am EDT on 7/9 to continue with the observation of RXTE J1550-564 (TOO). Radiation levels had fallen to acceptable levels and science operations were resumed.

3. Real time procedures.

In addition to the procedures above, a routine update of ephemeris coefficients was completed on 6/6, and the fifth Squeegee mode test was run on 6/9. This Squeege test, conducted at 7pm EDT on 6/9, tested an approach that combined on-chip summing, slower that normal parallel transfer time, and a 24 row exclude window.

Note that analysis indicates that the likely cause of the auto-unloading of momentum reported last week was a difference between the predicted and actual solar torque pressure at low Sun pitch angles, rather than higher than predicted gravity gradients at perigee.

The planned observations for the coming week are as follows.

------------------------------------------
PSS 1443+2724          ACIS-S          Jun 11
NGC 4472               ACIS-S
M82                    ACIS-I          Jun 12
Radiation Belts
Serpens X-1            ACIS-S/HETG     Jun 13
1812-12                ACIS-S
3C 273                 ACIS-S/LETG     Jun 14
3C 273                 ACIS-S/HETG
PSS 1317+3531          ACIS-S
PSS 1057+4555          ACIS-S
NGC 1399               ACIS-S/HETG
Radiation Belts                        Jun 15
Circinus Galaxy        ACIS-S/HETG
NGC 1399               ACIS-S/HETG     Jun 16
ARK 564                ACIS-S/HETG     Jun 17
Radiation Belts
TZ Crb                 ACIS-S/HETG     Jun 18

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports